OpenStudy (anonymous):

2*4 square root of 80

OpenStudy (mathstudent55):

This? \(2 \sqrt[4]{80}\)

OpenStudy (anonymous):

2*4*sqrt(80) 2*4*sqrt(16*5) 2*4*4*sqrt(5) 32*sqrt(5)

OpenStudy (mathstudent55):

\(2 \sqrt[4]{80}\) \(=2 \sqrt[4]{16 \times 5}\) \(=2 \sqrt[4]{16} \times \sqrt[4]{5}\) \(=2 \times 2 \times \sqrt[4]{5}\) \(=4 \sqrt[4]{5}\)

OpenStudy (anonymous):

@mathstudent55 is right then 2*(80)^1/4 2*(16*5)^1/4 2*2*(5)^1/4 \[4\sqrt[4]{5?}\]

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

Wlcm. Which one was it, square root or fourth root?

OpenStudy (anonymous):

it was 4 fourth root of 5